OpenStudy (anonymous):

The equation of motion of a particle is s = t^4 - 4 t, where s is in meters and t is in seconds. Assuming that t ge 0, answer the following questions Find the velocity v as a function of t. Answer: v(t) = 2. Find the acceleration a as a function of t. Answer: a(t) = 3. Find the acceleration after 2 seconds. Answer (in m/s^2): a(2) = 4. Find the acceleration when the velocity is 0. Answer (in m/s^2): a =

OpenStudy (kira_yamato):

v(t) = s'(t) = 4t^3 - 4 a(t) = v'(t) = 12t^2 a(2) = 12 (2^2) = 48 When v(t) = 0 t^3 = 1 t = 1 a(1) = 12 (1^2) = 12

OpenStudy (anonymous):

on the money thank you

OpenStudy (kira_yamato):

No problem ^^